OpenStudy (anonymous):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are -12 and 768, respectively. an = 3 • (-4)n + 1 an = 3 • 4n - 1 an = 3 • (-4)n - 1 an = 3 • 4n

OpenStudy (anonymous):

As we know Un=Up.r^n-p right?

OpenStudy (misssunshinexxoxo):

Which do you believe it is?

OpenStudy (anonymous):

honestly i have no clue how to do this

OpenStudy (misssunshinexxoxo):

T(n)=T(1)r^(n-1) => T(1)r=36--------(1) T(1)r^4=2304--------(2) (2)/(1)=> r^3=2304/36=> r=4 From (1), get T(1)*4=36=> T(1)=9 => the general term T(n)=9*4^(n-1)

OpenStudy (anonymous):

Hey are you sure you have written the choices correctly ?

OpenStudy (anonymous):

can i just plug them in and see which one gives me the 2nd as -12 and the 5th as 768

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Well yea thats the only way to solve but none works

OpenStudy (anonymous):

@misssunshinexxoxo which choice would it be?

OpenStudy (misssunshinexxoxo):

My best answer would be an = 3 • (-4)n - 1 should be good for this

OpenStudy (anonymous):

Wait r=-4 and not 4